evaluate (-16)^1/4
Answer by stanbon(75887)

evaluate (-16)^1/4
----------------
-16 = 16(cis(180)
---
(-16)^(1/4) = (16)^(1/4)(cis[(180+360k)/4] for k = 0,1,2,3
-----------------------
Ans: 4cis(45), 4cis(135),4cis(225),4cis(315)
=====================================================
Cheers,
Stan H.